Grade 12Discuss with Askiitians Tutors

3 identical dices are thrown simultaneously . what will b the probability of getting same no on the 3 dices?????????????

nivi  jain
14 Years agoGrade 12

3 Answers

Rahul Kumar
14 Years ago

Dear Nivi,

sample space = 63

and the number of favourable cases can be visualised as arranging 6 different objects in 3 different places 

this is same as the definition of permutation so the number of favourable cases is P(6,3)=6*5*4

so the probability = (6*5*4)/216 = 5/9

Ashwin Muralidharan IIT Madras
14 Years ago

Hi Nive,

 

Total Cases = 6*6*6 = 216

Favourable cases: (1,1,1) (2,2,2) (3,3,3)......(6,6,6)

Hence there are 6 favourable cases.

 

Probability = 6/216 = 1/36.

 

Hope it would help.

 

Regards,

Ashwin (IIT MadraS).

ayush jain
14 Years ago

hiiiiiiiiiiiiii.............

n(s)=6*6*6=216

n

n(e)=6

than probility =6/216=1/36